OpenStudy (anonymous):

Vectors Given OA=i+j, OB=4i-2j and OC=6i+kj, if A,B,C lie on a straight line,find the value of k. @phi

OpenStudy (phi):

can you find the direction vector form A to B? then write down the equation of the line that contains points A and B ?

OpenStudy (anonymous):

\[AB=3i-3j\]is it correct?

OpenStudy (phi):

yes, and we could simplify the *direction* to (1, -1) (we don't care about the length) and in this case, the arithmetic is easier. any way, what is the equation of the line ?

OpenStudy (anonymous):

\[AB=3(i-j)\]

OpenStudy (phi):

the equation is p0+ t* v where p0 is a point on the line (I would use A, but B also works) and v is the direction vector

OpenStudy (anonymous):

\[-i-j+3(i-j)\]

OpenStudy (phi):

we want the equation for any point OA=i+j or (1,1) vector from OA to OB is (3, -3) or 3(1, -1) we don't care about the 3 for a *direction*, so just use (1,-1) p= (1,1) + t (1, -1) follow so far?

OpenStudy (anonymous):

yes

OpenStudy (phi):

t=0 gives us (1,1) t=3 gives us (4, -2) now we want (1,1) + t(1, -1) = (6,k)

OpenStudy (phi):

if we use the first component, we can solve for t then use that t to find k

OpenStudy (anonymous):

1+t=6 t=5 1-t=k 1-5=k k=-4

OpenStudy (phi):

yes, point (5, -4) is on the line p= (1,1) + t (1, -1) (when t= 5)

OpenStudy (phi):

**point (6,-4)

OpenStudy (anonymous):

Thank you @phi :)